OpenStudy (anonymous):

Which of the following statements best describes the value of the expression 6x + 7 when x = 5? A. The result is a fraction. B. The result is a prime number. C. The result is a composite number. D. The result is a whole number that is neither prime nor composite.

OpenStudy (anonymous):

6x+7 when x=5 6(5)+7=30+7=37

OpenStudy (anonymous):

37 is ......... number?

OpenStudy (anonymous):